# A neural optimization framework for free-boundary diffeomorphic mapping problems and its applications

### v1

@@ Abstract
Free-boundary diffeomorphism optimization is a core ingredient in the surface mapping problem but remains notoriously difficult because the boundary is unconstrained and local bijectivity must be preserved under large deformation. Numerical Least-Squares Quasiconformal (LSQC) theory, with its provable existence, uniqueness, similarity-invariance and resolution-independence, offers an elegant mathematical remedy. However, the conventional numerical algorithm requires landmark conditioning, and cannot be applied into gradient-based optimization. We propose a neural surrogate, the Spectral Beltrami Network (SBN), that embeds LSQC energy into a multiscale mesh-spectral architecture. Next, we propose the SBN guided optimization framework SBN-Opt which optimizes free-boundary diffeomorphism for the problem, with local geometric distortion explicitly controllable. Extensive experiments on density-equalizing maps and inconsistent surface registration demonstrate our SBN-Opt’s superiority over traditional numerical algorithms.
@@ 1 Introduction
Diffeomorphic mappings are indispensable in computational geometry and medical imaging because they register surfaces while exactly preserving topology. Practical examples abound: population cartograms rescale countries to equalize densitychoi2018density; lyu2024bijective; surgeons “unfold” the cerebral cortex to a flat view for easier navigationbuchsbaum2012neural; and manufacturers flatten curved sheets to design cutting patterns for free-form partspietroni2022computational. In each case, the task is to map a given surface to the plane. Crucially, the shape of the planar image is often unknown in advance. The algorithm must infer the boundary shape so that the final map captures target properties such as area distribution, curvature, or feature alignment. We refer to this setting as the free-boundary diffeomorphism problem.
Established works like choi2015fast provide efficient algorithms for computing a conformal disk parameterization of an open surface, hence mathematically, the problem can be formulated as finding an optimal mapping on the unit disk $\mathbb{D}$ such that
@@
$f=\textbf{argmin}_{g:\mathbb{D}\rightarrow\mathbb{C}}E(g)\\ \text{ subject to }g\in\mathcal{C},$ (1)
where $E$ is a mapping energy functional and $\mathcal{C}$ denotes the admissible maps. Because practical applications demand bijectivity, we choose $\mathcal{C}$ as the space of diffeomorphisms.
The diffeomorphism optimization problem is generally challenging due to the nonlinearity and complexity of the constraint space. In the context of a fixed target domain, numerous effective methods have been developed for computing diffeomorphisms, each contributing valuable advancements to the field. For instance, flow-based techniques including LDDMM joshi2000landmark and SVF arsigny2006log, offer a principled way to generate diffeomorphisms by solving ordinary differential equations, which enables handling complex scenarios like large deformations. Other methods including smith2015bijective; fu2015computing; rabinovich2017scalable; lyu2024spherical; lyu2024bijective; kovalsky2015large, have also shown impressive results in practice. Nevertheless, each approach contains inherent trade-offs. For instance, flow-based strategies require solving an ODE to obtain the diffeomorphism, which can become computationally intensive, especially for high-resolution data. Some alternative methods rely on carefully tuning the optimization step size rabinovich2017scalable; lyu2024spherical; lyu2024bijective or post hoc projection kovalsky2015large to prevent local fold-over, or they introduce barrier energy terms smith2015bijective; fu2015computing to reduce the risk of folding. While such modifications are effective in many cases, they may not always guarantee bijectivity, and their effectiveness can depend sensitively on parameter choices and problem context.
Moreover, in a free-boundary diffeomorphism optimization problem, the shape of the target domain is not known in advance and typically depends on the optimization objective. This makes the problem significantly more challenging: the solver must simultaneously determine both the mapping and the geometry of the target domain. To address this challenge, most prior methods introduce additional constraints, such as fixing the positions of certain landmark points, adding artificial boundary frames, or prescribing a fixed outer boundary shape to anchor the solution and prevent degenerate outcomes qiu2020inconsistent; choi2018density. At the same time, many applications also require bounded geometric distortion. Balancing injectivity, distortion control, and optimizing target geometry produces a highly nonconvex optimization landscape, where algorithms can easily become trapped in undesirable, suboptimal solutions.
Therefore, our focus is developing a method for computing free-boundary diffeomorphisms, which has theoretic guarantee for bijectivity and can more efficiently control the local geometric distortion. To remove the need for boundary scaffolds while explicitly controlling local distortion, we recast the task as minimizing the least squares quasiconformal (LSQC) energy qiu2019computing
$E_{\text{LSQC}}(f,\mu)=\frac{1}{2}\int_{\mathbb{D}}\|f_{\bar{z}}-\mu f_{z}\|^{2}dA.$
As validated in Sec.5, numerical implementation of LSQC enjoys properties including existence and uniqueness of the minimizer, invariance due to a similarity transformation, and independence of mesh resolution. However, differentiating through the classical LSQC solver is computationally expensive, which limits its use in gradient-based optimization pipelines. To overcome this, we develop the Spectral Beltrami Network (SBN), a deep neural network designed to act as a neural surrogate for the LSQC solver that approximates the LSQC solution in a single fully differentiable forward pass. Given Beltrami coefficients $\mu$ (BCs) on the disk, SBN predicts a map that closely matches the output of the numerical solver without prescribing any boundary information. After the training of SBN stops, we freeze the parameters of SBN and treat BCs $\mu$ as the explicit, trainable input to SBN, shifting the optimization problem from searching directly in the unstable space of vertex coordinates to a well-structured space of BCs $\{\mu:\|\mu\|_{\infty}<1\}$ which essentially is the space of diffeomorphisms by the quasiconformal theory gardiner2000quasiconformal.
Within this, we propose a novel framework for the free-boundary diffeomorphism problem which solves the problem 1 by
@@
where $\mathcal{F}_{\theta^{*}}$ is the SBN, $g$ is the similarity transformation that $g(x,\phi,s,r)=se^{i\phi}x+r$, $E_{1}$ and $E_{2}$ are energy functionals on the mapping and BCs, respectively. Since the key component of the framework is SBN, we abbreviate the framework to SBN-Opt. The SBN-Opt changes from ’optimization over all deformations’ to ’optimization over diffeomorphisms’, giving us explicit, mathematically guaranteed control over both boundary freedom and local geometric distortion while maintaining the efficiency of a single forward pass.
The rest of the paper is organized as follows. Sec.2 reviews previous works related to our proposed framework. Sec.3 lists the contributions of this paper. Sec.4 introduces the mathematical theories involved in our work. In Sec.5 we propose and prove several key properties of LSQC energy. Then we introduce the architecture of Spectral Beltrami Network in Sec.6 and the optimization framework SBN-Opt in Sec.7 followed by implementation details in Sec.8 and extensive experiments in Sec.9. Lastly, a conclusion is given in Sec.10.
2 Previous works
Our research objective is to develop an efficient algorithm for the free-boundary diffeomorphism optimization problem, which can explicitly control the local geometric distortion. Therefore, we may review some representative works on the diffeomorphism optimization problem. Besides, since our strategy involves the quasiconformal theory, previous works related to computational quasiconformal geometry are also discussed in this section. Below we give a focused but not exhaustive overview of the most relevant prior works.
2.1 Diffeomorphism Optimization Problem
Diffeomorphic mapping provides one-to-one as well as smooth correspondences across different objects of interest, which serves as a powerful registration tool and has been successfully applied to many scenarios including image segmentation and image registration. Numerous works have been proposed for this problem. For example, the Large Diffeomorphic Distance Metric Mapping (LDDMM) joshi2000landmark maps the template domain onto the target one by defining and solving an ordinary differential equation (ODE). Stationary Vector Field (SVF) arsigny2006log is also defined by an ODE but different from LDDMM with time-varying speed flows, SVF provides stationary-speed vector fields and obtains the flow via exponential of the vector field. Later methods like DARTEL ashburner2007fast and diffeomorphic Demons vercauteren2009diffeomorphic further accelerate registration by exploiting Lie-algebra parameterizations, yet they still depend on careful regularization to avoid foldovers near steep deformations. To enforce injectivity more directly, barrier-energy terms smith2015bijective; fu2015computing that penalize near-zero Jacobian determinants and projection schemes kovalsky2015large that iteratively map invalid solutions back into the admissible set have been introduced, trading simplicity for additional optimization complexity and still only providing indirect control of distortion. Recent learning-based approaches krebs2021learning; amor2022resnet; wang2020deepflash predict velocity fields or full deformations in one forward pass, replacing costly iterative solvers, yet most did not involve contorl over local geometric distortion.
When comes to a free-boundary setting, achieving rigorous invertibility and tight distortion control without relying on boundary scaffolds or handcrafted regularizers remains challenging. One stream of methods is treating boundary points as individual optimization parameters and updating them in an iterative nonlinear numerical scheme. For example, Smith and Schaefer smith2015bijective designed a barrier distance function to avoid intersection of boundary edges. Another direction is to determine the boundary via the desired geometric properties in the inner region and geometry of the original surface. A typical work is the least squares conformal maps (LSCM) levy2002least, which obtain a free-boundary conformal parameterization of a surface via solving a linear system. An alternative ’sea-padding’ strategy choi2018density embeds the domain inside a large auxiliary rectangle whose rigid outer edge serves as an implicit boundary.
2.2 Computational Quasiconformal Geometry
Since the seminal introduction of discrete conformal mapping to computer graphics gu2002computing; gu2004genus; jin2008discrete; kharevych2006discrete, conformal techniques have become indispensable for surface processing. Lui and collaborators extended this framework to quasiconformal registration, developing the Linear Beltrami Solver (LBS) lui2013texture, which accommodates large deformations, guarantees bijectivity, and proved effective in medical imaging and surface matching lui2014teichmuller; choi2015flash; lyu2024spherical; lyu2024bijective. However, it still requires explicit boundary point correspondences. Drawing inspiration from LSCM levy2002least, Qiu and Lui introduced the least squares quasiconformal (LSQC) energy qiu2019computing, removing the need for prescribed boundary constraints and thus broadening applicability to tasks such as inconsistent surface registration qiu2020inconsistent; choi2020shape and partial welding zhu2022parallelizable. This boundary-free formulation underpins our work. More recently, several studies guo2023automatic; chen2024deep have coupled the quasiconformal theory with deep learning, training neural networks to map Beltrami coefficients to diffeomorphisms but these efforts simulate LBS within a fixed domain. Our Spectral Beltrami Network demonstrates, for the first time, that a neural surrogate can faithfully approximate the free-boundary LSQC solution, combining the advantage of deep learning with the theoretical guarantees of quasiconformal geometry.
3 Contributions
Our work has several key contributions and novel aspects to the field of free-boundary diffeomorphism optimization problem. Specifically, our contributions are as follows:
@@ 3 Contributions
    Several crucial properties of LSQC are proven, including existence and uniqueness, exact satisfaction of the Beltrami equation, guarantee of reconstructing diffeomorphism, invariance under similarity transformations, and resolution independence. These theoretical guarantees provide a solid mathematical foundation that enables its application to free-boundary diffeomorphism optimization.
@@
    A novel neural network architecture Spectral Beltrami Network (SBN) is designed to simulate LSQC, which innovatively combines:
    • Multiscale message-passing mechanisms that enable efficient information exchange across different mesh resolutions
    • Mesh Spectral Layers (MSL) that leverage Laplacian eigenmodes to capture global geometric dependencies
    As a differentiable surrogate of LSQC, SBN can compute free-boundary quasiconformal mappings on the unit disk, given Beltrami coefficients and two pinned points.
@@
    We established SBN-Opt, a novel optimization framework that uses the SBN as a differentiable engine to solve complex free-boundary diffeomorphism problems. Instead of merely training a network for one-shot predictions, SBN-Opt establishes a complete gradient-based optimization loop over the space of admissible BCs, equivalently the space of diffeomorphisms. Key features include:
    • Direct Optimization over Geometric Priors: The framework enables direct, gradient-based optimization of the underlying BCs $\mu$. This allows for explicit and intuitive control over local geometric properties like angle distortion.
    • Free of landmark conditions: Different from the traditional optimization scheme using the numerical LSQC algorithm which is landmark-dependent, SBN-Opt is designed to solve the optimization problem by simultaneously optimizing BCs and two pinned points thus is more flexible and able to applied to general free-boundary diffeomorphism problems.
    This framework effectively bridges the gap between the theoretical guarantees of quasiconformal geometry and the practical flexibility of deep learning, creating a powerful new tool for surface registration and parameterization.

### v2

@@ Abstract
Free-boundary diffeomorphism optimization, an important and widely occurring task in geometric modeling, computer graphics, and biological imaging, requires simultaneously determining a planar target domain and a locally bijective map with well-controlled distortion. We formulate this task through the least-squares quasiconformal (LSQC) operator and establish key structural properties of the LSQC minimizer, including well-posedness under mild conditions, invariance under similarity transformations, and resolution-independent behavior with stability under mesh refinement. We further analyze the sensitivity of the LSQC solution with respect to the Beltrami coefficient, establishing stability and differentiability properties that enable gradient-based optimization over the space of Beltrami coefficients. To make this differentiable formulation practical at scale and to facilitate the optimization process, we introduce the Spectral Beltrami Network (SBN), a multiscale mesh–spectral surrogate that approximates the LSQC solution operator in a single differentiable forward pass. This yields SBN-Opt, an optimization framework that searches over admissible Beltrami coefficients and pinning conditions to solve free-boundary diffeomorphism objectives with explicit distortion control. Extensive experiments on equiareal parameterization and inconsistent surface registration demonstrate consistent improvements over traditional numerical algorithms.
@@ 1 Introduction
Diffeomorphic mappings are indispensable in fields such as geometry processing, computer graphics, and medical imaging, where one must deform surfaces in a topology-preserving way so that measurements, annotations and features remain comparable after deformation. Practical examples abound in parameterization and registration problems: population cartograms rescale countries to equalize densityChoi and Rycroft (2018); Lyu et al. (2024a); surgeons “unfold” the cerebral cortex into a flat view for easier navigationBuchsbaum et al. (2012); and manufacturers flatten curved sheets to design cutting patterns for free-form partsPietroni et al. (2022); atlas-based registration and medical image segmentation require invertible deformations to align anatomical structures or biological featuresVercauteren et al. (2009); Yang et al. (2017). In many scenarios, the task involves finding the desired planar representation of an open surface. Crucially, the shape of the planar image is often needed to optimize. The algorithm must infer the boundary shape so that the final map captures target properties such as area distribution, curvature, or feature alignment. We refer to this setting as the free-boundary diffeomorphism problem.
Given a surface $\mathcal{M}$, a representative mathematical formulation is to find an optimal target domain $\Omega^{*}$ and an optimal mapping $f^{*}$ defined on $\mathcal{M}$:
@@
$(f^{*},\Omega^{*})=\textbf{argmin}_{\Omega,\;f:\mathcal{M}\rightarrow\Omega}E(f,\Omega)\quad\text{subject to }f\in\mathcal{C},$ (1)
where $E$ is a mapping objective and $\mathcal{C}$ denotes admissible maps. In applications, bijectivity is essential for the map. Beyond bijectivity, the map is usually required to have well-controlled geometric distortion so that it is physically plausible. Accordingly, the feasible set $\mathcal{C}$ should be understood as diffeomorphisms together with an explicit distortion constraint or regularization (e.g., bounds on local anisotropy), rather than arbitrary smooth bijections. This makes free-boundary diffeomorphism optimization challenging for two reasons: (i) bijectivity must be preserved but naive optimization over vertex coordinates or displacement fields is prone to foldovers; and (ii) distortion control must be enforced efficiently, ideally through a representation in which local distortion is explicit and can be constrained or penalized without projections or delicate step-size tuning. Therefore, our goal is to find a representation in which diffeomorphisms and their local distortion are encoded by an explicit quantity, and to optimize in that representation rather than in vertex coordinates. Specifically, we aim to identify a distortion-aware quantity that corresponds to diffeomorphisms, design algorithms to reconstruct the associated bijection from it, and then build an optimization framework directly over this quantity to control distortion efficiently and make free-boundary diffeomorphism optimization more stable and easier.
Related computational paradigms for diffeomorphic optimization.
Diffeomorphism optimization problems arise pervasively in surface parameterization and deformable registration, and the computational task can be viewed as optimizing over a (discretized) space of topology-preserving maps, ideally with (i) bijectivity and (ii) controlled distortion. Existing methods can be grouped into two broad paradigms, distinguished by what is taken as the optimization variable: (A) direct optimization of the map itself (e.g., vertex positions or displacement fields), and (B) optimization of a generating quantity from which a diffeomorphism is obtained by construction (e.g., velocity fields, momenta, or spline generators).
(A) Diffeomorphisms by directly optimizing the map.
In this paradigm, one treats the map $f$ as the primary unknown and then augments the task loss objective with regularizations or designs mechanisms intended to avoid local foldovers and reduce geometric distortion. In surface parameterization, a common pipeline first computes a bijective embedding and then improves the parameterization by minimizing a distortion functional subject to injectivity constraints Jiang et al. (2017); Su et al. (2020); Rabinovich et al. (2017); Smith and Schaefer (2015). The computational process from the initial embedding to the final 2D parameterization is a typical diffeomorphism optimization problem. Some methods first initialize using Tutte’s embedding or its variants and then optimize a flip-prevented distortion energy. For example, Smith et al. Smith and Schaefer (2015) combined a symmetric Dirichlet energy with a barrier preventing boundary self-intersection to reduce the distortion while ensuring no flips. Jiang et al. (2017); Su et al. (2020) added an auxiliary “scaffold” domain so that global non-overlap can be reduced to a local injectivity condition and also used a distortion metric or a barrier function. Most works need to impose distortion energies and design tailored numerical optimization strategies as well as mesh postprocessing methods. Same ideas can be seen in the deformable image registration research Liu et al. (2024); Jia et al. (2021); Balakrishnan et al. (2019). Many deep registration methods predict a dense displacement field in one forward pass and use smoothness regularization to discourage irregular deformations. VoxelMorph Balakrishnan et al. (2019) and VR-Net Jia et al. (2021) pursued unsupervised registration via a similarity term plus deformation regularization.
Optimizing the map is flexible and direct, but bijectivity and distortion control are usually achieved through auxiliary mechanisms (barriers, projections, penalty weights, line-search safeguards). This typically (i) yields sensitive hyperparameters, (ii) needs tailored optimization strategies and (iii) lacks direct control over geometric distortion (e.g., controlling local conformal distortion versus imposing energies on the singular values or Jacobian.)
(B) Diffeomorphisms by construction from generators.
In this paradigm, one does not optimize the map $f$ (or displacement field) directly. Instead, one optimizes a generating quantity from which a diffeomorphism is reconstructed, typically through integrating an ODE/PDE-defined flow or exponentiating a stationary velocity field. Classical representatives include LDDMM and related geodesic-shooting formulations Joshi and Miller (2000); Zhang et al. (2017); Beg et al. (2005), where the deformation is the flow of a velocity field; in practice the optimization can be posed either in terms of the velocity $v_{t}$ or in terms of its dual momentum variable. Demons-based methods Vercauteren et al. (2009, 2008) further reduced complexity by optimizing a stationary vector field $v$ and reconstructing the displacement field by squaring and scaling. Recent learning-based registration has adopted this generator viewpoint to obtain topology-preserving maps more reliably than direct displacement prediction. For example, Quicksilver Yang et al. (2017) predicted the momentum-parameterization of an LDDMM model in a patch-wise manner, aiming to approximate solutions of variational LDDMM optimization while retaining its diffeomorphic structure. Mok Mok and Chung (2020) proposed a fast symmetric diffeomorphic registration network that estimates forward and inverse transformations simultaneously, enabling fast computation of the inverse transformation. Han et al. Han et al. (2023) went a step further toward function-space parameterizations, which modeled a continuous velocity field using a neural network and then performed optimization in this generator space.
While generator-based parameterizations offer a principled route to diffeomorphisms and often improve robustness under large deformations, they typically regulate deformation through smoothness or kinetic-energy priors on the generator (velocity/momentum/spline coefficients). As a consequence, geometric distortion of the resulting map is controlled at most indirectly. This motivates seeking a representation in which the optimized quantity corresponds directly to (and quantitatively controls) geometric distortion, which we introduce next via quasiconformal geometry.
Quasiconformal (QC) geometry offers a principled representation of distortion-controlled diffeomorphisms via the Beltrami coefficient $\mu$, which encodes the conformality deviation and satisfies the correspondence between the set $\{\mu:\|\mu\|_{\infty}<1\}$ and the group of orientation-preserving quasiconformal maps. Discrete conformal and quasiconformal mappings have been central in surface processing Gu and Yau (2002); Gu et al. (2004); Kharevych et al. (2006); Lui et al. (2013). Lui et al. developed the Linear Beltrami Solver (LBS) Lui et al. (2013), which accommodates large deformations and guarantees bijectivity, and has been applied broadly in medical imaging and surface matching Lui et al. (2014); Choi et al. (2015); Lyu et al. (2024b, a). A key limitation is that LBS typically requires explicit boundary correspondences, a requirement that is not applicable in free-boundary settings. To remove prescribed boundary constraints, Qiu and Lui introduced the weighted least-squares quasiconformal energy Qiu et al. (2019), which allows a boundary-free formulation and has enabled tasks such as inconsistent surface registration Qiu and Lui (2020); Choi et al. (2020) and partial welding Zhu et al. (2022).
Methodology
Our idea is to seek a distortion-related representation of the mappings, which belongs to the second category of approaches. Given a planar domain $U$, We build on the $\ell^{2}$ norm of the Beltrami equation deviation and propose the unweighted least-squares quasiconformal energy (LSQC)
$E_{\text{LSQC}}(f,\mu)=\frac{1}{2}\int_{U}\|f_{\bar{z}}-\mu f_{z}\|^{2}\,dA.$ (2)
A key viewpoint in this paper is to treat LSQC not merely as an energy, but as a solution operator that maps Beltrami coefficients to free-boundary homeomorphisms. Concretely, as shown in the Section 3 for an admissible Beltrami coefficient $\mu$ with $\|\mu\|_{\infty}<1$, together with a minimal pinning condition (two pinned points to fix the similarity gauge and ensure non-triviality of the solution), LSQC yields a unique discrete map $f:U\to\mathbb{C}$ whose image domain $\Omega=f(U)$ is not prescribed but determined inherently by the Beltrami coefficient. Thus, we can view LSQC as an operator $\mathcal{F}$ that receives as input $\mu$$p_{i}$ in the mesh as well as their associated destinations $q_{i}\in\mathbb{C}$, and outputs a piecewise linear mapping $f$. In particular, given the Beltrami coefficient induced by a (piecewise linear) homeomorphism and two-point pinning, the LSQC exactly reconstructs the map. This leads to the following modeling insight:
Representation insight. Up to a similarity transformation, a free-boundary quasiconformal homeomorphism can be represented by its Beltrami coefficient $\mu$ together with the images of two pinned points.
We will establish that the LSQC solution map has the structure required for optimization over Beltrami coefficients.
  1. 1.
    Structure and well-posedness. Under mild conditions (including minimal pinning), the LSQC solution is well-defined and unique; it is invariant under similarity transforms and exhibits resolution-independent behavior that allows stable transfer across discretization levels.
  2. 2.
    Sensitivity. The LSQC minimizer depends stably on $\mu$ (in particular, it satisfies Lipschitz-type stability bounds under perturbations of $\mu$), and it is differentiable with respect to $\mu$ under standard regularity assumptions. These results justify gradient-based optimization over Beltrami coefficients.
Therefore, instead of optimizing directly over vertex coordinates (where injectivity is fragile), after obtaining a conformal parameterization $U=\phi(\mathcal{M})$ of the surface $\mathcal{M}$, we can reformulate the problem  (1) as an optimization problem over the structured Beltrami space,
$f^{*}=\textbf{argmin}_{\mu,p_{i}\in U.q_{i}\in\mathbb{C},f:U\rightarrow\mathbb{C}}E(f,\mu,p_{i},q_{i})\\ \text{ subject to }f,\mu\in\mathcal{C},$ (3)
where $\mathcal{C}$ is the space of admissible Beltrami coefficient such that $\|\mu\|_{\infty}<1,f=\mathcal{F}(\mu,p_{i},q_{i})$.
Although LSQC provides an elegant operator viewpoint, differentiating through a classical LSQC solver requires backpropagation through large sparse linear systems, which is computationally expensive for a high-resolution surface mesh. This computational barrier is one reason why many existing QC-based optimization pipelines update Beltrami variables through alternating or splitting strategies (often ADMM-like) Lui and Ng (2015); Lam and Lui (2014), rather than end-to-end gradients. To make the LSQC operator usable as a differentiable computational primitive at scale, we introduce the Spectral Beltrami Network (SBN), a multiscale mesh-spectral surrogate that approximates the free-boundary LSQC solution map in a single differentiable forward pass. After training, we freeze SBN and treat the Beltrami coefficient $\mu$ (together with pinning) as the explicit optimization variables, thereby shifting the optimization from unstable vertex-coordinate deformation spaces to the structured admissible set $\{\mu:\|\mu\|_{\infty}<1\}$, which corresponds to quasiconformal mappings in QC theory Gardiner and Lakic (2000). This yields SBN-Opt, an optimization framework that searches over admissible Beltrami coefficients and pinning conditions to solve free-boundary diffeomorphism objectives with explicit distortion control.
@@
Unlike prior QC optimization methods that incorporate $\mu$ through splitting/ADMM-style alternations, SBN-Opt enables gradient-based optimization directly in Beltrami coefficients through a differentiable surrogate of the LSQC operator. We validate the resulting theory-algorithm pipeline on tasks including equiareal parameterization and inconsistent surface registration.
Contributions
Our work provides a foundational framework for solving free-boundary diffeomorphism problems by re-characterizing them as optimization tasks within the space of Beltrami coefficients. Our contributions are summarized as follows:
@@
    Foundational Analysis of the LSQC Operator. We establish that the numerical Least-squares Quasiconformal (LSQC) energy defines a well-posed computational operator for free-boundary problems. Specifically, we prove that the discrete LSQC minimizer exists, is unique given at least two pinned points, and can exactly reconstruct the piecewise linear homeomorphism given its corresponding face-wise constant Beltrami fields. Furthermore, we prove that this operator possesses crucial geometric and computational invariances, including similarity-invariance and resolution-independence.
@@
    Theoretical Legitimacy of the Optimization Map. We provide the first rigorous analysis of the LSQC solution. We prove that the discrete LSQC minimizer depends Lipschitz-continuously on the input Beltrami field. Crucially, we establish the Fréchet differentiability of the mapping with respect to the Beltrami coefficients. These results demonstrate that gradient-based optimization in the space of Beltrami coefficient is mathematically sound and that small perturbations in the coefficients yield controlled, predictable changes in the resulting mapping.
@@
    The Spectral Beltrami Network (SBN) as a Differentiable Primitive. A major challenge in optimizing free-boundary diffeomorphisms is that backpropagating gradients through a numerical LSQC solver is computationally expensive and thus practically infeasible for iterative gradient-based pipelines. We introduce the SBN, a novel multiscale mesh-spectral architecture designed to act as a differentiable surrogate for the LSQC operator. By integrating multiscale message-passing with Mesh Spectral Layers (MSL), the SBN captures both local geometric interactions and global dependencies via transforming in the eigenspace of mesh Laplacian. This transforms the computationally expensive, implicitly-defined LSQC solver into a differentiable computational primitive capable of approximating the mapping in a single, fully-differentiable forward pass.
  4. 4.
[…]
  5. 5.
    Empirical validation on parameterization and registration tasks. We validate the full pipeline on equiareal parameterization and inconsistent surface registration. Extensive experiments show that SBN-Opt consistently produces mappings with significantly lower geometric distortion and higher accuracy than traditional numerical algorithms, meanwhile maintaining bijectivity and free-boundary characteristics.
Roadmap
The rest of the paper is organized as follows. Section 2, Section 3 and Section 4 develop the theoretical foundation. After introducing the required quasiconformal and discrete differential geometry background, we establish the structural properties of LSQC that make it suitable for optimization, including well-posedness under mild conditions, similarity-gauge invariance, and resolution-independent behavior, followed by a sensitivity analysis showing Lipschitz stability and Fréchet differentiability of the LSQC minimizer with respect to the Beltrami coefficient. Section 5 and Section 6 turn these results into a practical differentiable computational pipeline: we introduce the Spectral Beltrami Network (SBN) as a surrogate for the LSQC operator and then present SBN-Opt, an optimization framework over admissible Beltrami coefficients and pinning variables that enables explicit distortion control in free-boundary diffeomorphism objectives. Section 7 describes implementation details, and Section 8 reports experimental results on equiareal parameterization and inconsistent surface registration. Section 9 concludes with a discussion of achievements and future directions.